OpenStudy (anonymous):

Find all solutions if 0degrees is less than or equal to theta less than 360 degrees. tan^2 theta-tan theta=0

OpenStudy (campbell_st):

well factorise it by taking out the common factor \[\tan(\theta)(\tan(\theta) - 1) = 0\] the values that make the equation true are \[\tan(\theta) = 0...and... \tan(\theta) = 1\] you just need to find the angles that represent these values. Hope this make sense.

OpenStudy (anonymous):

thank you! would there be 4 solutions?

OpenStudy (campbell_st):

well 4 or 5... tan(x) = 0 x = 0, 180, 360

OpenStudy (anonymous):

i got 45, 225, 0, and 180. 360 cant be it because theta is less than 360. is that right?

OpenStudy (campbell_st):

ok... then thats a great step of answers